# KEK-TH-2773 Quantum vacuum energy and geometry of extra dimension

### v1

@@ B.1 Scalar sector
${\cal W}_{0}(y)$ $\equiv C_{0}(y;p_{\rm 4d}^{2})S_{0}^{\prime}(y;p_{\rm 4d}^{2})-S_{0}(y;p_{\rm 4d}^{2})C_{0}^{\prime}(y;p_{\rm 4d}^{2})=\frac{n^{4}(0)b(y)}{b(0)n^{4}(y)}.$ (B.2)
@@
${\cal W}_{L}(y)$ $\equiv C_{L}(y;p_{\rm 4d}^{2})S_{L}^{\prime}(y;p_{\rm 4d}^{2})-S_{L}(y;p_{\rm 4d}^{2})C_{L}^{\prime}(y;p_{\rm 4d}^{2})=\frac{n^{4}(L)b(y)}{b(L)n^{4}(y)}.$ (B.4)
@@
$S_{L}(y;p_{\rm 4d}^{2})$ $=\frac{b(0)n^{4}(L)}{n^{4}(0)b(L)}\left\{C_{0}(L;p_{\rm 4d}^{2})S_{0}(y;p_{\rm 4d}^{2})-S_{0}(L;p_{\rm 4d}^{2})C_{0}(y;p_{\rm 4d}^{2})\right\}.$ (B.5)
B.2 Spinor sector
@@
${\cal W}_{0}^{(\pm)}(y)$ $\equiv C_{0}^{(\pm)}(y;p_{\rm 4d}^{2})S_{0}^{(\pm)\prime}(y;p_{\rm 4d}^{2})-S_{0}^{(\pm)}(y;p_{\rm 4d}^{2})C_{0}^{(\pm)\prime}(y;p_{\rm 4d}^{2})=\frac{n(0)b(y)}{b(0)n(y)}.$ (B.8)
@@
${\cal W}_{L}^{(\pm)}(y)$ $\equiv C_{L}^{(\pm)}(y;p_{\rm 4d}^{2})S_{L}^{(\pm)\prime}(y;p_{\rm 4d}^{2})-S_{L}^{(\pm)}(y;p_{\rm 4d}^{2})C_{L}^{(\pm)\prime}(y;p_{\rm 4d}^{2})=\frac{n(L)b(y)}{b(L)n(y)}.$ (B.10)
B.3 Simple examples
@@
$S_{u}(y;p_{\rm 4d}^{2})$ $=S_{u}^{(\pm)}(y;p_{\rm 4d}^{2})=\frac{\sinh\left\{b_{\rm c}\sqrt{p_{\rm 4d}^{2}+M^{2}}(y-u)\right\}}{b_{\rm c}\sqrt{p_{\rm 4d}^{2}+M^{2}}},$ (B.12)
@@
$\hskip 62.59605pt\left.-I_{\nu}\left(e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right)K_{\nu}\left(e^{ky}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right)\right],$ (B.13)
@@
$B_{u}(p_{\rm 4d}^{2})$ $\equiv(2+\nu)I_{\nu}\left(e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right)+e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}I_{\nu+1}\left(e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right).$ (B.14)
@@
$\hskip 62.59605pt\left.-I_{\nu}\left(e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right)K_{\nu}\left(e^{ky}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right)\right\},$ (B.15)
@@
$B_{u}(p_{\rm 4d}^{2})$ $\equiv\left(\frac{1}{2}+\nu\right)I_{\nu}\left(e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right)+e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}I_{\nu+1}\left(e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right).$ (B.16)
Appendix C 5D propagator
C.1 Scalar sector
@@
$\langle 0|T\Phi(x^{\mu},y)\Phi(x^{\prime\nu},y^{\prime})|0\rangle$ $=\int\frac{d^{4}p}{i(2\pi)^{4}}\;e^{ip\cdot(x-x^{\prime})}\tilde{G}_{\rm B}(p_{\mu},y,y^{\prime}).$ (C.1)
@@
$=\frac{\delta(y-y^{\prime})}{a_{\rm sc}^{3}n^{4}(y)b(y)}.$ (C.2)
@@
$\tilde{G}_{\rm B}(p_{\mu},y,y^{\prime})$ $=\tilde{G}_{\rm B}(p_{\mu},y^{\prime},y).$ (C.3)
@@
$\partial_{y}\tilde{G}_{\rm B}(p_{\mu},0,y^{\prime})$ $=\partial_{y}\tilde{G}_{\rm B}(p_{\mu},L,y^{\prime})=0.$ (C.4)
By integrating (C.2) over infinitesimal interval that contain $y=y^{\prime}$, we obtain
$\lim_{y\to y^{\prime}}\left[\partial_{y}\tilde{G}_{\rm B}(p_{\mu},y,y^{\prime})-\partial_{y}\tilde{G}_{\rm B}(p_{\mu},y,y^{\prime})\right]$ $=-\frac{b(y^{\prime})}{n^{4}(y^{\prime})}.$ (C.5)
From (C.4), we can express $\tilde{G}_{B}(p_{\mu},y,y^{\prime})$ in terms of the basis functions in Appendix B.1 as
$\tilde{G}_{\rm B}(p_{\mu},y,y^{\prime})$ $=\vartheta(y-y^{\prime})C_{L}(y;p_{\rm 4d}^{2})\alpha_{>}(y^{\prime};p_{\rm 4d}^{2})+\vartheta(y^{\prime}-y)C_{0}(y;p_{\rm 4d}^{2})\alpha_{<}(y^{\prime};p_{\rm 4d}^{2}),$ (C.6)
where $\vartheta(y)$ is the Heaviside step function. By using (C.3) and (C.5), we find that
@@
$\alpha_{<}(y^{\prime};p_{\rm 4d}^{2})$ $=-\frac{b(y^{\prime})C_{L}(y^{\prime};p_{\rm 4d}^{2})}{a_{\rm sc}^{3}n^{4}(y^{\prime})\left(C_{0}C_{L}^{\prime}-C_{0}^{\prime}C_{L}\right)(y^{\prime};p_{\rm 4d}^{2})}=\frac{b(L)C_{L}(y^{\prime};p_{\rm 4d}^{2})}{a_{\rm sc}^{3}n^{4}(L)C_{0}^{\prime}(L;p_{\rm 4d}^{2})}.$ (C.7)
@@
$\left(C_{0}C_{L}^{\prime}-C_{0}^{\prime}C_{L}\right)(y^{\prime};p_{\rm 4d}^{2})$ $=-C_{0}^{\prime}(L;p_{\rm 4d}^{2})\frac{n^{4}(L)b(y^{\prime})}{b(L)n^{4}(y^{\prime})},$ (C.8)
which follows from (B.2) and (B.5). Note that $\tilde{G}_{\rm B}(p_{\mu},y,y^{\prime})$ is a function of $p_{\rm 4d}^{2}$. Thus it is denoted as $\tilde{G}_{\rm B}(p_{\rm 4d}^{2},y,y^{\prime})$ in the main text.
C.2 Spinor sector
@@
$=\frac{\delta(y-y^{\prime})}{a_{\rm sc}^{3}b(y)}\mbox{$1$}_{4},$ (C.11)
@@
$\hat{G}_{\rm F}(p_{\mu},y,y^{\prime})$ $\equiv-\begin{pmatrix}\hat{G}_{\chi\lambda}(p_{\mu},y,y^{\prime})&\hat{G}_{\chi\chi}(p_{\mu},y,y^{\prime})\\ \hat{G}_{\lambda\lambda}(p_{\mu},y,y^{\prime})&\hat{G}_{\lambda\chi}(p_{\mu},y,y^{\prime})\end{pmatrix}$. (C.12)
@@
$\hat{G}_{\lambda\chi}^{\dagger}(p_{\mu},y,y^{\prime})$ $=\hat{G}_{\chi\lambda}(p_{\mu},y^{\prime},y).$ (C.13)
@@
$\left.\left(\partial_{y}+bM\right)\hat{G}_{\lambda\chi}\right|_{y=0}$ $=\left.\left(\partial_{y}+bM\right)\hat{G}_{\lambda\chi}\right|_{y=L}=0.$ (C.15)
@@
$\hskip 14.22636pt\times\left\{\vartheta(y-y^{\prime})H_{L\lambda}(y;p_{\rm 4d}^{2})H_{0\lambda}(y^{\prime};p_{\rm 4d}^{2})+\vartheta(y^{\prime}-y)H_{0\lambda}(y;p_{\rm 4d}^{2})H_{L\lambda}(y^{\prime};p_{\rm 4d}^{2})\right\},$ (C.16)
@@
$H_{L\lambda}(y;p_{\rm 4d}^{2})$ $\equiv C_{L}^{(+)}(y;p_{\rm 4d}^{2})-b(L)MS_{L}^{(+)}(y;p_{\rm 4d}^{2}).$ (C.17)
The basis functions $C_{0,L}^{(\pm)}(y;p_{\rm 4d}^{2})$ and $S_{0,L}^{(\pm)}(y;p_{\rm 4d}^{2})$ are defined in Appendix B.2.
Appendix D Expansions for large momentum
@@
D.1 Scalar sector
D.1.1 Basis functions
@@
$\left[\partial_{y}^{2}+\left(\frac{4n^{\prime}}{n}-\frac{b^{\prime}}{b}\right)\partial_{y}-\frac{b^{2}}{n^{2}}\rho^{2}-b^{2}M^{2}\right]f_{u}(y;\rho)$ $=0.$ (D.1)
@@
$f_{u}(y;\rho)$ $=\frac{\exp\left[\rho\,{\cal U}_{u}(y)\right]}{n^{3/2}(y)}\left[1+\frac{g_{u,1}(y)}{\rho}+\frac{g_{u,2}(y)}{\rho^{2}}+\frac{g_{u,3}(y)}{\rho^{3}}+\cdots\right],$ (D.2)
@@
${\cal U}_{u}(y)$ $\equiv\int_{u}^{y}d\tilde{y}\;\frac{b(\tilde{y})}{n(\tilde{y})}$, (D.3)
@@
$g_{u,m\geq 2}^{\prime}$ $=-\frac{n}{2b}\left[g_{u,m-1}^{\prime\prime}+\left(\frac{n^{\prime}}{n}-\frac{b^{\prime}}{b}\right)g^{\prime}_{u,m-1}-\left(\frac{3n^{\prime\prime}}{2n}+\frac{9n^{\prime 2}}{4n^{2}}-\frac{3n^{\prime}b^{\prime}}{2nb}+b^{2}M^{2}\right)g_{u,m-1}\right].$ (D.4)
Note that when $f_{u}(y;\rho)$ is a solution of (D.1), $f_{u}(y;-\rho)$ is also a solution of it. Hence we can express the basis functions as linear combinations of them. Since the basis functions are functions of $\rho^{2}$, they should be proportional to $f_{u}(y;\rho)\pm f_{u}(y;-\rho)$. Taking into account the conditions in (B.1) and (B.3), the basis function $C_{u}(y;\rho^{2})$ is found to be
$C_{u}(y;\rho^{2})$ $=\frac{n^{3/2}(u)}{2}\left[f_{u}(y;\rho)+f_{u}(y;-\rho)\right],$ (D.5)
@@
$g_{2l+1}(u)$ $=-\frac{n(u)}{b(u)}g_{2l}^{\prime}(u)+\frac{3n^{\prime}(u)}{2b(u)}g_{2l}(u)=-\frac{n(u)}{b(u)}g_{2l}^{\prime}(u).\;\;\;\;\;\left(l=1,2,3,\cdots\right)$ (D.6)
@@
$S_{u}(y;\rho^{2})$ $=\frac{n^{5/2}(u)}{2\rho b(u)}\left[f_{u}(y;\rho)-f_{u}(y;-\rho)\right],$ (D.7)
@@
$g_{2l-1}(u)$ $=0,\;\;\;\;\;g_{2l}(u)=-\frac{n(u)}{b(u)}g_{2l-1}^{\prime}(u).\;\;\;\;\;\left(l=1,2,3,\cdots\right)$ (D.8)
D.1.2 5D propagator
@@
${\cal U}_{0}(y)$ $=\int_{0}^{y}d\tilde{y}\;\frac{b(\tilde{y})}{n(\tilde{y})}\geq 0,\;\;\;\;\;{\cal U}_{L}(y^{\prime})=\int_{L}^{y^{\prime}}d\tilde{y}\;\frac{b(\tilde{y})}{n(\tilde{y})}\leq 0.$ (D.9)
@@
$C_{0}^{\prime}(L;\rho^{2})$ $=\frac{\rho n^{3/2}(0)b(L)}{2n^{5/2}(L)}e^{\rho{\cal U}_{0}(L)}\left[1+\frac{{\cal G}_{1}}{\rho}+\frac{{\cal G}_{2}}{\rho^{2}}+\frac{{\cal G}_{3}}{\rho^{3}}+\cdots\right],$ (D.10)
@@
$\simeq\frac{e^{-\rho\Delta(y,y^{\prime})}}{2\rho a_{\rm sc}^{3}n^{3/2}(y)n^{3/2}(y^{\prime})}\left[1+\frac{{\cal H}_{1}(y,y^{\prime})}{\rho}+\frac{{\cal H}_{2}(y,y^{\prime})}{\rho^{2}}+\frac{{\cal H}_{3}(y,y^{\prime})}{\rho^{3}}+\cdots\right],$ (D.12)
@@
$\vdots$ (D.13)
From (D.12), we also have
$\partial_{y}\partial_{y}^{\prime}\tilde{G}_{\rm B<}(\rho^{2},y,y^{\prime})$ $\simeq\frac{e^{-\rho\Delta(y,y^{\prime})}}{2a_{\rm sc}^{3}n^{3/2}(y)n^{3/2}(y^{\prime})}\left[\rho{\cal S}_{-1}+{\cal S}_{0}+\frac{{\cal S}_{1}}{\rho}+\frac{{\cal S}_{2}}{\rho^{2}}+\frac{{\cal S}_{3}}{\rho^{3}}+\cdots\right],$ (D.14)
@@
$\quad+\left\{\partial_{y}\partial_{y}^{\prime}-\frac{3n^{\prime}(y^{\prime})}{2n(y^{\prime})}\partial_{y}-\frac{3n^{\prime}(y)}{2n(y)}\partial_{y}^{\prime}+\frac{9n^{\prime}(y)n^{\prime}(y^{\prime})}{4n(y)n(y^{\prime})}\right\}{\cal H}_{l-1}(y,y^{\prime}),$ (D.15)
@@
${\cal H}_{0}(y,y^{\prime})$ $\equiv 1,\;\;\;\;\;{\cal H}_{l}(y,y^{\prime})\equiv 0.\;\;\;\;\;\left(l<0\right)$ (D.16)
D.2 Spinor sector
D.2.1 Basis functions

### v2

@@ C.1 Scalar sector
${\cal W}_{0}(y)$ $\equiv C_{0}(y;p_{\rm 4d}^{2})S_{0}^{\prime}(y;p_{\rm 4d}^{2})-S_{0}(y;p_{\rm 4d}^{2})C_{0}^{\prime}(y;p_{\rm 4d}^{2})=\frac{n^{4}(0)b(y)}{b(0)n^{4}(y)}.$ (C.2)
@@
${\cal W}_{L}(y)$ $\equiv C_{L}(y;p_{\rm 4d}^{2})S_{L}^{\prime}(y;p_{\rm 4d}^{2})-S_{L}(y;p_{\rm 4d}^{2})C_{L}^{\prime}(y;p_{\rm 4d}^{2})=\frac{n^{4}(L)b(y)}{b(L)n^{4}(y)}.$ (C.4)
@@
$S_{L}(y;p_{\rm 4d}^{2})$ $=\frac{b(0)n^{4}(L)}{n^{4}(0)b(L)}\left\{C_{0}(L;p_{\rm 4d}^{2})S_{0}(y;p_{\rm 4d}^{2})-S_{0}(L;p_{\rm 4d}^{2})C_{0}(y;p_{\rm 4d}^{2})\right\}.$ (C.5)
C.2 Spinor sector
@@
${\cal W}_{0}^{(\pm)}(y)$ $\equiv C_{0}^{(\pm)}(y;p_{\rm 4d}^{2})S_{0}^{(\pm)\prime}(y;p_{\rm 4d}^{2})-S_{0}^{(\pm)}(y;p_{\rm 4d}^{2})C_{0}^{(\pm)\prime}(y;p_{\rm 4d}^{2})=\frac{n(0)b(y)}{b(0)n(y)}.$ (C.8)
@@
${\cal W}_{L}^{(\pm)}(y)$ $\equiv C_{L}^{(\pm)}(y;p_{\rm 4d}^{2})S_{L}^{(\pm)\prime}(y;p_{\rm 4d}^{2})-S_{L}^{(\pm)}(y;p_{\rm 4d}^{2})C_{L}^{(\pm)\prime}(y;p_{\rm 4d}^{2})=\frac{n(L)b(y)}{b(L)n(y)}.$ (C.10)
C.3 Simple examples
@@
$S_{u}(y;p_{\rm 4d}^{2})$ $=S_{u}^{(\pm)}(y;p_{\rm 4d}^{2})=\frac{\sinh\left\{b_{\rm c}\sqrt{p_{\rm 4d}^{2}+M^{2}}(y-u)\right\}}{b_{\rm c}\sqrt{p_{\rm 4d}^{2}+M^{2}}},$ (C.12)
@@
$\hskip 62.59605pt\left.-I_{\nu}\left(e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right)K_{\nu}\left(e^{ky}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right)\right],$ (C.13)
@@
$B_{u}(p_{\rm 4d}^{2})$ $\equiv(2+\nu)I_{\nu}\left(e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right)+e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}I_{\nu+1}\left(e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right).$ (C.14)
@@
$\hskip 62.59605pt\left.-I_{\nu}\left(e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right)K_{\nu}\left(e^{ky}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right)\right\},$ (C.15)
@@
$B_{u}(p_{\rm 4d}^{2})$ $\equiv\left(\frac{1}{2}+\nu\right)I_{\nu}\left(e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right)+e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}I_{\nu+1}\left(e^{ku}\frac{b_{\rm c}\sqrt{p_{\rm 4d}^{2}}}{k}\right).$ (C.16)
Appendix D 5D propagator
D.1 Scalar sector
@@
$\langle 0|T\Phi(x^{\mu},y)\Phi(x^{\prime\nu},y^{\prime})|0\rangle$ $=\int\frac{d^{4}p}{i(2\pi)^{4}}\;e^{ip\cdot(x-x^{\prime})}\tilde{G}_{\rm B}(p_{\mu},y,y^{\prime}).$ (D.1)
@@
$=\frac{\delta(y-y^{\prime})}{a_{\rm sc}^{3}n^{4}(y)b(y)}.$ (D.2)
@@
$\tilde{G}_{\rm B}(p_{\mu},y,y^{\prime})$ $=\tilde{G}_{\rm B}(p_{\mu},y^{\prime},y).$ (D.3)
@@
$\partial_{y}\tilde{G}_{\rm B}(p_{\mu},0,y^{\prime})$ $=\partial_{y}\tilde{G}_{\rm B}(p_{\mu},L,y^{\prime})=0.$ (D.4)
By integrating (D.2) over infinitesimal interval that contain $y=y^{\prime}$, we obtain
$\lim_{y\to y^{\prime}}\left[\partial_{y}\tilde{G}_{\rm B}(p_{\mu},y,y^{\prime})-\partial_{y}\tilde{G}_{\rm B}(p_{\mu},y,y^{\prime})\right]$ $=-\frac{b(y^{\prime})}{a_{\rm sc}^{3}n^{4}(y^{\prime})}.$ (D.5)
From (D.4), we can express $\tilde{G}_{B}(p_{\mu},y,y^{\prime})$ in terms of the basis functions in Appendix C.1 as
$\tilde{G}_{\rm B}(p_{\mu},y,y^{\prime})$ $=\vartheta(y-y^{\prime})C_{L}(y;p_{\rm 4d}^{2})\alpha_{>}(y^{\prime};p_{\rm 4d}^{2})+\vartheta(y^{\prime}-y)C_{0}(y;p_{\rm 4d}^{2})\alpha_{<}(y^{\prime};p_{\rm 4d}^{2}),$ (D.6)
where $\vartheta(y)$ is the Heaviside step function. By using (D.3) and (D.5), we find that
@@
$\alpha_{<}(y^{\prime};p_{\rm 4d}^{2})$ $=-\frac{b(y^{\prime})C_{L}(y^{\prime};p_{\rm 4d}^{2})}{a_{\rm sc}^{3}n^{4}(y^{\prime})\left(C_{0}C_{L}^{\prime}-C_{0}^{\prime}C_{L}\right)(y^{\prime};p_{\rm 4d}^{2})}=\frac{b(L)C_{L}(y^{\prime};p_{\rm 4d}^{2})}{a_{\rm sc}^{3}n^{4}(L)C_{0}^{\prime}(L;p_{\rm 4d}^{2})}.$ (D.7)
@@
$\left(C_{0}C_{L}^{\prime}-C_{0}^{\prime}C_{L}\right)(y^{\prime};p_{\rm 4d}^{2})$ $=-C_{0}^{\prime}(L;p_{\rm 4d}^{2})\frac{n^{4}(L)b(y^{\prime})}{b(L)n^{4}(y^{\prime})},$ (D.8)
which follows from (C.2) and (C.5). Note that $\tilde{G}_{\rm B}(p_{\mu},y,y^{\prime})$ is a function of $p_{\rm 4d}^{2}$. Thus it is denoted as $\tilde{G}_{\rm B}(p_{\rm 4d}^{2},y,y^{\prime})$ in the main text.
D.2 Spinor sector
@@
$=\frac{\delta(y-y^{\prime})}{a_{\rm sc}^{3}b(y)}\mbox{$1$}_{4},$ (D.11)
@@
$\hat{G}_{\rm F}(p_{\mu},y,y^{\prime})$ $\equiv-\begin{pmatrix}\hat{G}_{\chi\lambda}(p_{\mu},y,y^{\prime})&\hat{G}_{\chi\chi}(p_{\mu},y,y^{\prime})\\ \hat{G}_{\lambda\lambda}(p_{\mu},y,y^{\prime})&\hat{G}_{\lambda\chi}(p_{\mu},y,y^{\prime})\end{pmatrix}$. (D.12)
@@
$\hat{G}_{\lambda\chi}^{\dagger}(p_{\mu},y,y^{\prime})$ $=\hat{G}_{\chi\lambda}(p_{\mu},y^{\prime},y).$ (D.13)
@@
$\left.\left(\partial_{y}+bM\right)\hat{G}_{\lambda\chi}\right|_{y=0}$ $=\left.\left(\partial_{y}+bM\right)\hat{G}_{\lambda\chi}\right|_{y=L}=0.$ (D.15)
@@
$\hskip 14.22636pt\times\left\{\vartheta(y-y^{\prime})H_{L\lambda}(y;p_{\rm 4d}^{2})H_{0\lambda}(y^{\prime};p_{\rm 4d}^{2})+\vartheta(y^{\prime}-y)H_{0\lambda}(y;p_{\rm 4d}^{2})H_{L\lambda}(y^{\prime};p_{\rm 4d}^{2})\right\},$ (D.16)
@@
$H_{L\lambda}(y;p_{\rm 4d}^{2})$ $\equiv C_{L}^{(+)}(y;p_{\rm 4d}^{2})-b(L)MS_{L}^{(+)}(y;p_{\rm 4d}^{2}).$ (D.17)
The basis functions $C_{0,L}^{(\pm)}(y;p_{\rm 4d}^{2})$ and $S_{0,L}^{(\pm)}(y;p_{\rm 4d}^{2})$ are defined in Appendix C.2.
Appendix E Expansions for large momentum
@@
E.1 Scalar sector
E.1.1 Basis functions
@@
$\left[\partial_{y}^{2}+\left(\frac{4n^{\prime}}{n}-\frac{b^{\prime}}{b}\right)\partial_{y}-\frac{b^{2}}{n^{2}}\rho^{2}-b^{2}M^{2}\right]f_{u}(y;\rho)$ $=0.$ (E.1)
@@
$f_{u}(y;\rho)$ $=\frac{\exp\left[\rho\,{\cal U}_{u}(y)\right]}{n^{3/2}(y)}\left[1+\frac{g_{u,1}(y)}{\rho}+\frac{g_{u,2}(y)}{\rho^{2}}+\frac{g_{u,3}(y)}{\rho^{3}}+\cdots\right],$ (E.2)
@@
${\cal U}_{u}(y)$ $\equiv\int_{u}^{y}d\tilde{y}\;\frac{b(\tilde{y})}{n(\tilde{y})}$, (E.3)
@@
$g_{u,m\geq 2}^{\prime}$ $=-\frac{n}{2b}\left[g_{u,m-1}^{\prime\prime}+\left(\frac{n^{\prime}}{n}-\frac{b^{\prime}}{b}\right)g^{\prime}_{u,m-1}-\left(\frac{3n^{\prime\prime}}{2n}+\frac{9n^{\prime 2}}{4n^{2}}-\frac{3n^{\prime}b^{\prime}}{2nb}+b^{2}M^{2}\right)g_{u,m-1}\right].$ (E.4)
Note that when $f_{u}(y;\rho)$ is a solution of (E.1), $f_{u}(y;-\rho)$ is also a solution of it. Hence we can express the basis functions as linear combinations of them. Since the basis functions are functions of $\rho^{2}$, they should be proportional to $f_{u}(y;\rho)\pm f_{u}(y;-\rho)$. Taking into account the conditions in (C.1) and (C.3), the basis function $C_{u}(y;\rho^{2})$ is found to be
$C_{u}(y;\rho^{2})$ $=\frac{n^{3/2}(u)}{2}\left[f_{u}(y;\rho)+f_{u}(y;-\rho)\right],$ (E.5)
@@
$g_{2l+1}(u)$ $=-\frac{n(u)}{b(u)}g_{2l}^{\prime}(u)+\frac{3n^{\prime}(u)}{2b(u)}g_{2l}(u)=-\frac{n(u)}{b(u)}g_{2l}^{\prime}(u).\;\;\;\;\;\left(l=1,2,3,\cdots\right)$ (E.6)
@@
$S_{u}(y;\rho^{2})$ $=\frac{n^{5/2}(u)}{2\rho b(u)}\left[f_{u}(y;\rho)-f_{u}(y;-\rho)\right],$ (E.7)
@@
$g_{2l-1}(u)$ $=0,\;\;\;\;\;g_{2l}(u)=-\frac{n(u)}{b(u)}g_{2l-1}^{\prime}(u).\;\;\;\;\;\left(l=1,2,3,\cdots\right)$ (E.8)
E.1.2 5D propagator
@@
${\cal U}_{0}(y)$ $=\int_{0}^{y}d\tilde{y}\;\frac{b(\tilde{y})}{n(\tilde{y})}\geq 0,\;\;\;\;\;{\cal U}_{L}(y^{\prime})=\int_{L}^{y^{\prime}}d\tilde{y}\;\frac{b(\tilde{y})}{n(\tilde{y})}\leq 0.$ (E.9)
@@
$C_{0}^{\prime}(L;\rho^{2})$ $=\frac{\rho n^{3/2}(0)b(L)}{2n^{5/2}(L)}e^{\rho{\cal U}_{0}(L)}\left[1+\frac{{\cal G}_{1}}{\rho}+\frac{{\cal G}_{2}}{\rho^{2}}+\frac{{\cal G}_{3}}{\rho^{3}}+\cdots\right],$ (E.10)
@@
$\simeq\frac{e^{-\rho\Delta(y,y^{\prime})}}{2\rho a_{\rm sc}^{3}n^{3/2}(y)n^{3/2}(y^{\prime})}\left[1+\frac{{\cal H}_{1}(y,y^{\prime})}{\rho}+\frac{{\cal H}_{2}(y,y^{\prime})}{\rho^{2}}+\frac{{\cal H}_{3}(y,y^{\prime})}{\rho^{3}}+\cdots\right],$ (E.12)
@@
$\vdots$ (E.13)
From (E.12), we also have
$\partial_{y}\partial_{y}^{\prime}\tilde{G}_{\rm B<}(\rho^{2},y,y^{\prime})$ $\simeq\frac{e^{-\rho\Delta(y,y^{\prime})}}{2a_{\rm sc}^{3}n^{3/2}(y)n^{3/2}(y^{\prime})}\left[\rho{\cal S}_{-1}+{\cal S}_{0}+\frac{{\cal S}_{1}}{\rho}+\frac{{\cal S}_{2}}{\rho^{2}}+\frac{{\cal S}_{3}}{\rho^{3}}+\cdots\right],$ (E.14)
@@
$\quad+\left\{\partial_{y}\partial_{y}^{\prime}-\frac{3n^{\prime}(y^{\prime})}{2n(y^{\prime})}\partial_{y}-\frac{3n^{\prime}(y)}{2n(y)}\partial_{y}^{\prime}+\frac{9n^{\prime}(y)n^{\prime}(y^{\prime})}{4n(y)n(y^{\prime})}\right\}{\cal H}_{l-1}(y,y^{\prime}),$ (E.15)
@@
${\cal H}_{0}(y,y^{\prime})$ $\equiv 1,\;\;\;\;\;{\cal H}_{l}(y,y^{\prime})\equiv 0.\;\;\;\;\;\left(l<0\right)$ (E.16)
E.2 Spinor sector
E.2.1 Basis functions